Imagine that you are a participant in a game show, where "free" money is given away. There are 4 prizes hidden on a game board with 16 spaces. One prize is worth $4000, another is worth $1500, and two are worth $1000 each.
If you select one of the prize spaces, you will get the related prize. However, if you select any of the other spaces, you will have to pay $50, as penalty for not making the ‘wise’ choice
In this game show, you are actually given a choice:
• Choice #1: You are offered a sure prize of $400 in cash. You can take the money and leave.
• Choice #2: Take your chance and play the game.
What would be your choice? Take the money and run, or play the game? Please explain your decision.

Answer by stanbon(57203)

Imagine that you are a participant in a game show, where "free" money is given away. There are 4 prizes hidden on a game board with 16 spaces. One prize is worth $4000, another is worth $1500, and two are worth $1000 each.
If you select one of the prize spaces, you will get the related prize. However, if you select any of the other spaces, you will have to pay $50, as penalty for not making the ‘wise’ choice
In this game show, you are actually given a choice:
• Choice #1: You are offered a sure prize of $400 in cash. You can take the money and leave.
• Choice #2: Take your chance and play the game.
What would be your choice? Take the money and run, or play the game? Please explain your decision.
--------------------
Random "winning" values: 5000,1500,1000,1000
Corresponding probabilities are all 4/16 = 1/4
---------------------
Expected "Winnings": (5000 + 1500 + 1000 + 1000 - 50*12)/16 = $493.75
------
Since the expected winnings is not that much better than $400,
take the $400 and quit the game.
Cheers,
Stan H.